OpenStudy (anonymous):

Ari has a total of 22 coins consisting of pennies and nickels. The total value of the coins is $0.54. How many of each type of coin does Ari have? _______ pennies and ________ nickels

OpenStudy (anonymous):

30 nickles 9 pennies

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

yeah buts its funny

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

8 nickels 14 pennies

OpenStudy (anonymous):

and thats right its not a joke

OpenStudy (anonymous):

lol

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{n+p=22\\~~~~\\ 5n+p=54\\~~~~\\ \text{solve this}}\end{align}\)

OpenStudy (anonymous):

8 and 14